# Metric Graph Kernels via the Tropical Torelli Map

### v1

@@ A.2 Proof of Theorem 1
A.3 Example
(a) Example graph
(b) Orientation and minimal spanning tree
Figure 6: Illustration of Algorithm 1. On the right panel, the orientation is indicated by arrow and the minimal spanning tree is colored in red.
We demonstrate the computation of the tropical Torelli map on a toy example shown in Figure 6(a). The graph $G$ has 6 nodes and 8 edges. The length of each edge is defined as $\ell(e_{i})=i$. We compute the minimal spanning tree as $T=\{e_{1},e_{2},e_{4},e_{5},e_{6}\}$, which is shown in Figure 6(b). After fixing an orientation, each edge not in $T$ determines a 1-cycle in the following:
$\sigma_{1}$ $=e_{3}-e_{2}+e_{1},$
$\sigma_{2}$ $=e_{7}-e_{5}-e_{6},$
$\sigma_{3}$ $=e_{8}-e_{6}-e_{4}-e_{2}.$
Thus the cycle-edge incidence matrix is given by
The matrix of the tropical Torelli map can be computed by
$Q=MLM^{\top}=\begin{bmatrix}6&0&2\\ 0&18&6\\ 2&6&20\end{bmatrix}.$
Appendix B Graph Kernels
B.1 Backgrounds of Kernels
Kernels form a foundational class of similarity functions in machine learning and data analysis, which are often used to capture nonlinear relationships by embedding data into high-dimensional feature spaces. Let $\mathcal{X}$ be a set and $k:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ be a symmetric function. Commonly $k$ is required to be positive definite in the sense that for any finite set of points $x_{1},x_{2},\ldots,x_{n}\in\mathcal{X}$ the matrix $K_{ij}=k(x_{i},x_{j})$ is positive semi-definite, i.e., for any real coefficients $c_{1},c_{2},\ldots,c_{n}$, we have
$\sum_{i=1}^{n}\sum_{j=1}^{n}c_{i}c_{j}k(x_{i},x_{j})\geq 0.$
For any positive definite kernel $k$, there exists a unique Hilbert space $\mathcal{H}$ known as the reproducing kernel Hilbert space (RKHS), along with a map $\phi:\mathcal{X}\to\mathcal{H}$ called the feature map, such that the value of the kernel is given by the inner product of the feature map [37]
$k(x,y)=\langle\phi(x),\phi(y)\rangle_{\mathcal{H}}.$
A typical construction of kernel is through the distance on a space. Let $d(x,y)$ be a distance function on $\mathcal{X}$, the distance is said to be conditionally negative type if for any finite set of points $x_{1},x_{2},\ldots,x_{n}\in\mathcal{X}$ and any real coefficients $c_{1},c_{2},\ldots,c_{n}$ such that $\sum_{i=1}^{n}c_{i}=0$, we have
$\sum_{i=1}^{n}\sum_{j=1}^{n}c_{i}c_{j}d(x_{i},x_{j})\leq 0.$
If $d(x,y)$ is conditionally negative type, then the kernel defined by $k(x,y)=\exp(-\gamma d(x,y)^{2})$ is positive definite for any $\gamma>0$, which is known as Schoenberg’s theorem [27]. The Euclidean distance is a typical example of conditionally negative type distance. However, the 2-Wasserstein distance is not of conditionally negative type [24], and thus the induced kernel is indefinite. In practice indefinite kernels are also widely used, and the corresponding learning methodology is called indefinite learning. In general, an indefinite kernel corresponds to a Reproducing Kernel Kreĭn Space (RKKS), on which many RKHS-based techniques can be extended [22, 34].
B.2 Compared Graph Kernels
@@
We list the graph kernels used for comparison in the benchmark dataset experiments below:
@@
  • The Edge Histogram (EH) kernel. The EH kernel compares graphs based on the histogram of discrete edge weights or labels [30]. It is the most computationally efficient graph kernel. However it cannot capture the graph topology.
@@
  • The Graphlet Sampling (GS) kernel. The GS kernel measures the similarity between graphs via counting matching graphlets, i.e., small subgraphs with $k$ vertices [28]. Since enumerating all graphlets is computationally expensive, we set the maximal size of a graphlet to be $k=6$, and the maximal number of samples to be 300.
@@
  • The ODD-STh (OS) kernel The OS kernel compares graphs based on the Directed Acyclic Graph (DAG) decomposition [11]. For each node in a graph, a breadth-first search (BFS) tree or DAG is constructed up to a fixed depth, capturing the local neighborhood structure in a rooted format. For unlabeled (or constantly labeled) graphs, the OS kernel is able to capture the structural differences.
@@
  • The Shortest Path (SP) kernel. The SP kernel compares graphs based on the shortest path lengths between all pairs of nodes [5]. The original SP kernel uses a base kernel function on the node labels at the path endpoints and the path lengths, which is computationally expensive. We use the simplified version using the histogram of shortest paths, which is also the implemented version in GraKeL.

### v2

@@ Appendix C Technical Proofs
Finally, for any graph refinement $G^{\prime}\geq G$, we can decompose it as $G^{\prime}=G_{k}\geq G_{k-1}\geq\ldots\geq G_{0}=G$ where each $G_{i+1}\geq G_{i}$ is given by one edge subdivision. Due to the invariance under edge subdivision we have $Q(G^{\prime})=Q(G)$, as desired. ∎
C.2 Derivation of the Bures–Wasserstein Distance
We now provide a brief proof of the following formula:
$\min_{U_{1},U_{2}\in O(g)}\|U_{1}\sqrt{Q_{1}}-U_{2}\sqrt{Q_{2}}\|_{F}^{2}=\mathrm{Tr}(Q_{1})+\mathrm{Tr}(Q_{2})-2\mathrm{Tr}\bigg(\sqrt{\sqrt{Q_{1}}Q_{2}\sqrt{Q_{1}}}\bigg)$ (5)
for positive semi-definite real matrices $Q_{1},Q_{2}\in\mathrm{PSD}(g)$. A general proof for positive semi-definite Hermitian matrices can be found in Bhatia et al. (2019).
Proof.
First, notice that
$\min_{U_{1},U_{2}\in O(g)}\|U_{1}\sqrt{Q_{1}}-U_{2}\sqrt{Q_{2}}\|_{F}^{2}$ $=\min_{U\in O(g)}\|U\sqrt{Q_{1}}-\sqrt{Q_{2}}\|_{F}^{2}$
$=\min_{U\in O(g)}\mathrm{Tr}\big(\big(U\sqrt{Q_{1}}-\sqrt{Q_{2}}\big)^{\top}\big(U\sqrt{Q_{1}}-\sqrt{Q_{2}}\big)\big)$
$=\min_{U\in O(g)}\mathrm{Tr}(Q_{1})+\mathrm{Tr}(Q_{2})-\mathrm{Tr}\big(\sqrt{Q_{2}}\sqrt{Q_{1}}U^{\top}\big)-\mathrm{Tr}\big(\sqrt{Q_{1}}\sqrt{Q_{2}}U\big)$
$=\mathrm{Tr}(Q_{1})+\mathrm{Tr}(Q_{2})-2\max_{U\in O(g)}\mathrm{Tr}\big(\sqrt{Q_{1}}\sqrt{Q_{2}}U\big).$
Now consider the SVD decomposition $\sqrt{Q_{1}}\sqrt{Q_{2}}=W\Lambda V^{\top}$. We have
$\mathrm{Tr}\big(\sqrt{Q_{1}}\sqrt{Q_{2}}U\big)=\mathrm{Tr}(W\Lambda V^{\top}U)=\mathrm{Tr}(\Lambda V^{\top}UW)=\sum_{i=1}^{g}\lambda_{i}\theta_{i},$
where $\theta_{i}$’s are the diagonal elements of $V^{\top}UW\in O(g)$. Since $|\theta_{i}|\leq 1$, we have $\sum_{i=1}^{g}\lambda_{i}\theta_{i}\leq\sum_{i=1}^{g}\lambda_{i}=\mathrm{Tr}(\Lambda)$, where equality holds if and only if $V^{\top}UW=I$. Since the singular values of $\sqrt{Q_{1}}\sqrt{Q_{2}}$ are the square roots of the eigenvalues of $(\sqrt{Q_{1}}\sqrt{Q_{2}})(\sqrt{Q_{1}}\sqrt{Q_{2}})^{\top}=\sqrt{Q_{1}}Q_{2}\sqrt{Q_{1}}$, we have
$\mathrm{Tr}(\Lambda)=\mathrm{Tr}\bigg(\sqrt{\sqrt{Q_{1}}Q_{2}\sqrt{Q_{1}}}\bigg),$
which is the last term of (5). ∎
C.3 Proof of Theorem 4.2
[…]
Definition C.1.
Let $G$ be a graph equipped with length function $\ell:E(G)\to\mathbb{R}_{+}$. A 3-edge-connectivization of $(G,\ell)$ is a new weighted graph $(G^{3},\ell^{3})$ such that
  • $G^{3}$ is obtained by contracting all bridges of $G$, and for each C1-set $S$ of $G$, contracting all but one edge of $S$; and
  • For any $e\in E(G^{3})$, let $S_{e}$ be the corresponding C1-set of $e$ in $G$, then $\ell^{3}(e)=\displaystyle\sum_{e^{\prime}\in S_{e}}\ell(e^{\prime})$.
Defintion C.1 is a simplified version of 3-edge-connectivization of a tropical curve $\Gamma$ (Brannetti et al., 2011, Lemma-Definition 5.3.2). The following injectivity result for the tropical Torelli map applies directly to our setting.
Theorem C.2.
(Brannetti et al., 2011, Theorem 5.3.3) Let $\Gamma$ and $\Gamma^{\prime}$ be two tropical curves of genus $g$. Then $\mathrm{Jac}(\Gamma)\cong\mathrm{Jac}(\Gamma^{\prime})$ if and only if $\Gamma^{3}$ is 2-isomorphic to $\Gamma^{\prime 3}$.
We now proceed to our proof of Theorem 4.2.
Proof.
Assume $(G,\ell)$ is 2-isomorphic to $(G^{\prime},\ell^{\prime})$, and let $\phi:E(G)\to E(G^{\prime})$ be a 2-isomorphism. Since $\ell$ is generic, it determines a unique minimal spanning forest $\mathrm{SF}_{G}$ for $G$. Similarly, $\ell^{\prime}$ determines a unique minimal spanning forest $\mathrm{SF}_{G^{\prime}}$ for $G^{\prime}$. Since $\phi$ preserves edge lengths, it sends $\mathrm{SF}_{G}$ to $\mathrm{SF}_{G^{\prime}}$ and preserves the order of non-tree edges in $G$ and $G^{\prime}$. By construction we have $Q(G)=Q(G^{\prime})$ and $k_{\mathrm{TTW}}(G,G^{\prime})=1$.
[…]
C.4 Proof of Theorem 4.3
Proof.
By Theorem 3.2, for a weighted graph $G$ of genus $g$ equipped with a generic length function, the map $\mathcal{M}_{\leq g}\to\mathrm{PSD}(g),\,G\mapsto Q(G)$ is well-defined. For any $G_{1},\ldots,G_{n}\in\mathcal{M}_{\leq g}$ and real numbers $c_{1},\ldots,c_{n}\in\mathbb{R}$, since the Gaussian kernel $k(x,y)=\exp(-\gamma\|x-y\|^{2})$ is known to be positive definite on $\mathbb{R}^{g^{2}}$, we have
$\sum_{i,j=1}^{n}c_{i}c_{j}\exp\big(-\gamma\|Q(G_{i})-Q(G_{j})\|_{F}^{2}\big)\geq 0,$
which implies that $k_{\mathrm{TTE}}$ is a positive definite kernel on $\mathcal{M}_{\leq g}$. ∎
Appendix D Additional Numerical Verifications
In this section we present additional numerical results in Section 5.1. Specifically, we test practical computational runtime of our metric graph kernels. We further investigate the stability of kernel matrices under random subsampling, test their positive semi-definiteness, and analyze their sensitivity to perturbations in edge lengths.
D.1 Verifying Computation Runtime.
We ran experiments on synthetic datasets to verify the computation runtime on our proposed classes of sparsity based on genus. The details are described in Section 5.1. Figure 6 shows that the runtime to compute the TTE kernel is in alignment with the growth of the graph genus, while the runtime to compute the TTW kernel grows faster with graph density.
D.2 Subsampling Stability
We provide experimental verifications of our subsampling method described in Section 4.1 to map high-dimensional SPD matrices to the same dimension by extracting a random principal submatrix.
As discussed in Section 4.1, it is necessary to consider the reduction loss in terms of the Frobenius norm since each matrix element contains topological information of the underlying metric graph. Suppose the original SPD matrix $Q$ is of size $g$ and the target size is $g_{0}$. Let $I\subseteq\{1,2,\ldots,g\}$ denote the index set of size $g_{0}$. Then in principle, we consider the following optimization problem
$\max_{\begin{subarray}{c}|I|=g_{0}\\ I\subseteq\{1,2,\ldots,g\}\end{subarray}}\sum_{i,j\in I}|Q_{ij}|^{2}.$ (6)
However, an exact solution of (6) can only be obtained by ranging over all index sets of size $g_{0}$, whose complexity is $\binom{g}{g_{0}}$. Therefore, instead of solving an exact solution, we simply subsample one index set $I$ randomly.
We test the stability of the TTW and TTE kernel matrices with respect to random matrix reduction using the following experimental setting: We generate a set of $N=20$ graphs with genus $g\in\{100,200,300\}$. For each graph $G$, we compute its tropical Torelli matrices $Q(G)$ and randomly extract a principal submatrix $\widehat{Q}_{\alpha}(G)$ of size $g_{0}=[\alpha g]$, where the ratio has range $\alpha\in\{0,1,0.2,\ldots,0.9\}$. Then we compute the kernel matrix $K$ using original $Q(G_{i})$, and the kernel matrix $\widehat{K}$ using reduced $\widehat{Q}_{\alpha}(G_{i})$. We repeat the above computations for $B=10$ times and compute the following mean relative error:
$\frac{1}{B}\sum_{i=1}^{B}\frac{\|\widehat{K}-K\|_{F}^{2}}{\|K\|_{F}^{2}}.$
The mean relative error and standard deviation are displayed in Table 3 and Table 4, from which we see that both kernels are stable with respect to random principal submatrix selection. We also see that the TTW kernel is more stable than the TTE kernel since it has smaller mean relative error for every reduction ratio.
Table 3: Numerical stability of random principal submatrix selection for the TTW kernel.
$g=100$ 0.870(0.003) 0.756( 0.004) 0.651(0.003) 0.551(0.005) 0.452(0.004) 0.355(0.003) 0.263(0.004) 0.170(0.003) 0.088(0.003)
$g=200$ 0.874(0.002) 0.760(0.003) 0.659(0.003) 0.555(0.002) 0.461(0.001) 0.365(0.004) 0.273(0.003) 0.181(0.003) 0.090(0.002)
$g=300$ 0.875(0.002) 0.766(0.002) 0.664(0.002) 0.562(0.001) 0.466(0.002) 0.372(0.003) 0.277(0.002) 0.183(0.002) 0.093(0.001)
Table 4: Numerical stability of random principal submatrix selection for the TTE kernel.
$g=100$ 0.951(0.004) 0.916(0.006) 0.854(0.009) 0.771(0.014) 0.673(0.020) 0.548(0.023) 0.437(0.022) 0.309(0.015) 0.188(0.025)
$g=200$ 0.970(0.003) 0.932(0.008) 0.876(0.006) 0.793(0.011) 0.696(0.012) 0.600(0.009) 0.474(0.016) 0.335(0.023) 0.171(0.012)
$g=300$ 0.972(0.001) 0.936(0.003) 0.864(0.004) 0.799(0.012) 0.698(0.010) 0.612(0.012) 0.474(0.016) 0.348(0.020) 0.161(0.013)
D.3 Kernel Positive Definiteness
Recall that a kernel function $k$ is positive definite if and only if for any $x_{1},\ldots,x_{N}\in\mathcal{X}$, the kernel matrix $k(x_{i},x_{j})]_{N\times N}$ is positive semi-definite (Section B.2). To test the positive definiteness of TTW and TTE kernels, we examine the smallest nonzero eigenvalues of their corresponding kernel matrices on different sets of data. Here we set a threshold $\mathrm{tol}=10^{-6}$ and an eigenvalue $\lambda$ is considered as nonzero if $|\lambda|>\mathrm{tol}$.
First we test how the smallest nonzero eigenvalue of the kernel matrices varies with the genus bound $g_{0}$. Specifically, for a fixed number of graphs $N\in\{30,50,70\}$, we generate $N$ random weighted graphs of genus $g\leq g_{0}$, where $g_{0}$ ranges from $20$ to $125$. For each configuration, we compute the corresponding kernel matrix and its smallest nonzero eigenvalue. The experiment is repeated $B=20$ times for each value of $N$. The mean and standard deviation of the smallest nonzero eigenvalues are presented in Figure 7. The results show that, for each fixed $N$, the smallest nonzero eigenvalues of both kernel matrices tend to increase with respect to the genus bound $g_{0}$. Moreover, for large genus bounds, the smallest nonzero eigenvalues of both kernel matrices remain positive.
(a) TTW
(b) TTE
Figure 7: Smallest nonzero eigenvalues of kernel matrices v.s. graph genus ($N\in\{30,50,70\}$)
We then test how the smallest nonzero eigenvalue of the kernel matrices varies with the number of graphs $N$. Similarly, for a fixed genus bound $g_{0}\in\{20,40,60\}$, we generate $N$ random weighted graphs of genus $g\leq g_{0}$, where $N$ ranges from $50$ to $190$. The experiment is repeated $B=20$ times for each value of $g_{0}$. The mean and standard deviation of the smallest nonzero eigenvalues are presented in Figure 8. The results show that, for each fixed $g_{0}$, the smallest nonzero eigenvalues of both kernel matrices tend to decrease with respect to the number of graphs $N$ which is also the size of the kernel matrix. For the TTE kernel, the smallest nonzero eigenvalue is always positive; for the TTW kernel, we observe small negative eigenvalues when $g_{0}$ is small and $N$ is large. However, we do not observe significant instability of the TTW kernel.
(a) TTW
(b) TTE
Figure 8: Smallest nonzero eigenvalues of kernel matrices v.s. number of graphs ($g_{0}\in\{20,40,60\}$)
D.4 Stability Under Edge Length Perturbations
We investigate the robustness of the TTW and TTE kernels with respect to edge length perturbations. Specifically, for a given graph $G$, we assume an additive noise model applied independently to each edge
$\ell(e_{i})=\bar{\ell}(e_{i})+\epsilon_{i},\quad\epsilon_{i}\sim\mathrm{Unif}(0,R).$ (7)
In general, suppose that the edge length matrix is perturbed by $L+\Delta L$. Then, as shown in (1) in the construction of the tropical Torelli matrix, we have
$\Delta Q=M_{T}\Delta L_{T}M^{\top}_{T}+\Delta L_{g}.$
Therefore the error satisfies $\|\Delta Q\|\leq\|M_{T}\|^{2}\|\Delta L_{T}\|+\|\Delta L_{g}\|$, which is proportional to the noise magnitude $R$ under (7).
To evaluate the variation of kernel matrices with respect to edge length perturbations, we adopt the following experimental setting: for fixed number of graphs $N$ and fixed genus bound $g_{0}$, we randomly sample $N$ graphs with genus no greater than $g_{0}$. Then we compute the true kernel matrix $K_{\mathrm{true}}$. For the same set of weighted graphs, we add noise to every edge according to (7), with noise magnitude $R$ ranging from $0.01$ to $5.12$. After obtaining the noisy kernel matrix $K_{\mathrm{noisy}}$, we compute the relative error $\|K_{\mathrm{noisy}}-K_{\mathrm{true}}\|^{2}_{F}/\|K_{\mathrm{true}}\|^{2}_{F}$.
Figure 9 shows the log-log plot of relative errors with respect to noise magnitudes for fixed $g_{0}=30$ and $N\in\{60,70,100\}$; Figure 10 shows the corresponding log-log plot for fixed $N=30$ and $g_{0}\in\{20,40,60\}$. From both figures, we observe that the TTW and TTE kernels are robust to edge-length perturbations.
(a) TTW
(b) TTE
Figure 9: Relative error of kernel matrices with respect to edge length perturbations ($N\in\{60,70,100\}$)
(a) TTW
(b) TTE
Figure 10: Relative error of kernel matrices with respect to edge length perturbations ($g_{0}\in\{20,40,60\}$)
Appendix E Additional Experimental Results on Benchmark Datasets
E.1 Compared Graph Kernels
@@
We list the graph kernels used for comparison in the benchmark dataset experiments, explain their core ideas, and describe the hyperparameters that require tuning.
@@
  • The Edge Histogram (EH) kernel. The EH kernel compares graphs based on the histogram of discrete edge weights or labels (Sugiyama and Borgwardt, 2015). It is the most computationally efficient graph kernel. However it cannot capture the graph topology. The EH kernel does not have any hyperparameters.
@@
  • The Graphlet Sampling (GS) kernel. The GS kernel measures the similarity between graphs via counting matching graphlets, i.e., small subgraphs with $k$ vertices (Shervashidze et al., 2009). Since enumerating all graphlets is computationally expensive, the counting and matching steps are commonly approximated by sampling a specified number of graphlets up to a given size. The hyperparameters for GS kernel are (1) maximal size of a graphlet, which is commonly set as $k=3\sim 6$; (2) the maximal number of samples which we set to be $B=300\sim 500$.
@@
  • The ODD-STh (OS) kernel The OS kernel compares graphs based on the Directed Acyclic Graph (DAG) decomposition (Da San Martino et al., 2012). For each node in a graph, a breadth-first search (BFS) tree or DAG is constructed up to a fixed depth, capturing the local neighborhood structure in a rooted format. For unlabeled (or constantly labeled) graphs, the OS kernel is able to capture the structural differences. The hyperparameter for the OS kernel is the maximal depth of the subtree pattern, which is commonly set to be $h=2\sim 5$.
@@
  • The Shortest Path (SP) kernel. The SP kernel compares graphs based on the shortest path lengths between all pairs of nodes (Borgwardt and Kriegel, 2005). The original SP kernel uses a base kernel function on the node labels at the path endpoints and the path lengths, which is computationally expensive. We use the simplified version using the histogram of shortest paths, which is also the implemented version in GraKeL. The SP kernel does not have any hyperparameters.
  • The Weisfeiler–Lehman (WL) Kernel. The WL kernel is a fast and powerful graph kernel based on the Weisfeiler–Lehman graph isomorphism test algorithm (Shervashidze et al., 2011). Its core idea is to iteratively relabel each node by aggregating the multiset of labels of its neighbors, thus capturing local neighborhood structures. In particular, we will use the WL subtree kernel implemented in GraKeL which is widely used in the literature. The hyperparameter for the WL kernel is the number of refinement iterations, which is commonly set to be $h=5\sim 10$.